Shift report 20000823 08:00-16:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Wed Aug 23 2000 - 16:53:34 EDT

  • Next message: Konstantin Olchanski: "BRAHMS DAQ run control updated"

    ------------------------------------------------------------
    Date: Wed Aug 23 16:53:33 2000, Shift: 08:00-16:00
    Supervisor: R. Debbe, with: Claus Ekman
    Log book#4 pages 146 - 150
    
    Shift Summary:
    ------------------------------------------------------------
    Machine status:  One store started at 8:15 with ZDC coincidence rates around
    120 Hz   we collected data in run 2410 and then decided to scale
    down trigger 4 ( by a factor of 2) to reduce dead time.
    
    2 Achievements:  Runs 2410 and 2411 with some 250K trigger 1
    events. MRS at 90 2Kgauss pol B FFS at 5 with T1 off
    
    At 11:00 a "quench link event " dumped the beam and RHIC went into
    maintenance mode and we were able to enter the IR to work on T1
    
    T1 readout board was removed and inspected for broken wires;
    another broken anode wire was found.  The new broken wires
    was located over the same guard strip and is the nearest neighbor
    to the wire removed on 8/21.
    
    The chances that this second wire broke because of electrostatic
    instabilities is high, we thus decided that the new readout
    board should be repaired in the lab. Meanwhile we will use
    the old board.
    
    During test in the lab the old board performed well but not
    as well as the new one. We decided that it would be good to
    have the ability to put some positive voltage on the field
    shaping wires, in order to have that choice we had to 
    disconnect the gating grid pulser because it was using the 
    cable intended for the field shapers.  We will start by
    taking data at 8 degrees where we know that the gating grid 
    was not necessary.
    
     
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Wed Aug 23 2000 - 16:56:49 EDT